# Hidden heavy flavor tetraquarks in the Born-Oppenheimer approximation

### v1

@@ Abstract
The Born-Oppenheimer approximation is one of the very successful tools for solving the hydrogen atom problem. The experimental discovery of hidden heavy flavor tetraquarks, $Q\bar{Q}q\bar{q}$ ($Q=c,b$ and $q=u,d,s$), provides great possibilities for the hydrogen-bond-like structure of the Quantum Chromodynamics version. In this work, considering that the colors of $Q\bar{Q}$ and $q\bar{q}$ are both 8, the color-spin wave functions of the tetraquark formed by the color coupling $8\otimes 8\rightarrow 1$ are constructed. Then we use the Born-Oppenheimer approximation to calculate the mass spectra of hidden heavy flavor tetraquarks and evaluate their spatial sizes. The results show that some of the hidden heavy flavor exotic hadrons discovered experimentally can be well explained as this type of hydrogen bond tetraquark structure. In addition, some candidates for tetraquark bound states formed by color-spin interactions are predicted and may be compact tetraquark states.
@@ I Introduction
Quantum chromodynamics (QCD) provides a rich framework for understanding hadrons, extending beyond the conventional mesons (one quark, one antiquark) and baryons (three quarks) to encompass exotic configurations such as multiquark states, hybrids, and glueballs Godfrey and Isgur (1985); Klempt and Zaitsev (2007); Lebed et al. (2017). Among the multiquark states, the study of tetraquark states (composed of two quarks and two antiquarks) have received significant attention Jaffe (1977a, b). Notably, certain hidden heavy flavor tetraquarks, $Q\bar{Q}q\bar{q}$ ($Q=c,b$ and $q=u,d,s$), can be interpreted as analogues of hydrogen bonds within QCD version. This allows the interaction between heavy quark-antiquark pairs and light quarks can be treated using the Born-Oppenheimer (BO) approximation Esposito et al. (2013); Braaten et al. (2014); Maiani et al. (2019a, b); Mutuk (2024); Berwein et al. (2024, 2024); Germani et al. (2025). This approach decouples the heavy quark dynamics from the light quark dynamics, enabling systematic studies of tetraquark spectra and structure, based on the fact that the heavy quarks are much heavier than the light quarks.
@@
Recently, there is increasing experimental evidence proving the existence of possible hidden heavy tetraquark states. In 2020, the first candidate for a charged hidden-charm tetraquark state with strangeness, $Z_{cs}(3985)^{-}$, was discovered in the processes of $e^{+}e^{-}\rightarrow K^{+}D_{s}^{-}D^{\ast 0}$ and $K^{+}D_{s}^{\ast-}D^{0}$ by the BESIII collaboration Ablikim et al. (2021). Its non-zero charge characteristic enables it to be clearly distinguished from hybrid or charmonium. Theoretically, $Z_{cs}(3985)^{-}$ is generally interpreted as tetraquark state $cs\bar{c}\bar{u}$ with a quantum number assignment $J^{P}=1^{+}$ Meng et al. (2020); Karliner and Rosner (2021); Giron et al. (2021); Guo et al. (2023); Liu et al. (2025). Later in 2021 the LHCb collaboration reported two hidden-charm exotic structures in the decay $B^{+}\rightarrow J/\psi\phi K^{+}$, referred to as $Z_{cs}(4000)^{+}$ and $Z_{cs}(4220)^{+}$ Aaij et al. (2021). Their quark composition is thought to be $cu\bar{c}\bar{s}$, and their spin-parity quantum quantum numbers are determined to be $J^{P}=1^{+}$ and $1^{+}$ (or $1^{-}$), respectively Aaij et al. (2021). However, theoretical studies suggest that the $Z_{cs}(4000)^{+}$ and $Z_{cs}(4220)^{+}$ may be candidates for charmonia A. and Dhir (2024); Bokade and Bhaghyesh (2025), tetraquarks Li et al. (2024); Vogt (2024); Yu et al. (2024), or molecules Yang et al. (2021); Chen (2024). It is worth noting that the BESIII collaboration recently observed a new $Z_{cs}(4123)^{-}$ in the processes of $e^{+}e^{-}\rightarrow K^{+}D_{s}^{\ast-}D^{\ast 0}+c.c.^{\ast}$, and its quantum number assignment may be $J^{P}=1^{+}$ Ablikim et al. (2023). The discovery of the $Z_{cs}(4123)^{-}$ may add a new member to the family of charged hidden-charm tetraquark with strangeness. In 2023, A near-threshold peaking structure $X(3960)$ is observed in the $D_{s}^{+}D_{s}^{-}$ invariant mass spectrum by the LHCb collaboration, and the quantum number assignment $J^{PC}=1^{++}$ is favored Aaij et al. (2023). Research shows that the nature of $X(3960)$ may also be a tetraquark $cs\bar{c}\bar{s}$ Badalian and Simonov (2023); Hoffer et al. (2024); Chen et al. (2025). Obviously, these exotic hadrons all have the common components $Q\bar{Q}q\bar{q}$, and figuring out their nature will help us better understand QCD.
@@
In this article, we apply the BO approximation to calculate the mass spectra and spatial sizes of tetraquark $Q\bar{Q}q\bar{q}$. The heavy $Q\bar{Q}$ forms a color octet, which combines with a light $q\bar{q}$ octet to produce a color-singlet tetraquark, that is, the color coupling is represented as $8\otimes 8\rightarrow 1$. The light quark (or antiquark) interactions are treated as perturbations to the dominant heavy-quark potential, allowing the Schrödinger equation for the heavy quark separation to be solved numerically. Given the same amount of momentum, the heavy quarks move much more slowly than the light quarks. The calculation of diquark-antidiquark paradigm was extended to tetraquarks. Using $X(3872)$ as input, other possible tetraquark states were studied.
@@
This paper is organized as follows. In Sec. II, We give a brief introduction of the tetraquark Hamiltonian based on the BO approximation. We obtain the BO potential using the perturbation method and construct the basis vector wave functions of the tetraquark system $Q\bar{Q}q\bar{q}$. In Sec. III, we determine the model parameters, such as the masses of quarks and the color coupling strengths. The mass spectra and spatial sizes of various hidden heavy tetraquark states are calculated and discussed in this section. We summarize in Sec. IV. The nature units $c=\hbar=1$ are used throughout.
@@ II.1 Hamiltonian in the Born-Oppenheimer approximation
$\begin{split}H=&\sum_{i=1}^{4}\left(\frac{p_{i}^{2}}{2m_{i}}\right)+V(x_{A,B})\\ &+V_{I}(x_{A,B},x_{1,2})+V_{\text{conf}},\end{split}$ (1)
where the coordinates and masses of the heavy quarks are denoted as $x_{A,B}$ and $M$, the coordinates and masses of the light quarks are $x_{1,2}$ and $m$. $\sum_{i=1}^{4}\left(\frac{p_{i}^{2}}{2m_{i}}\right)$ represents the kinetic energy of the quarks. $V(x_{A,B})$ describes the interaction between the heavy quarks. $V_{I}(x_{A,B},x_{1,2})$ describes the interactions between the heavy quarks, light quarks, and the light quarks themselves. $V_{\text{conf}}$ describes the color confinement in the singlet state of the hidden-flavor tetraquark system.
In the quark model, the spatial degree of freedom can be encoded in the model parameters. the Hamiltonian for the $S$-wave tetraquark can be written in a simplified form as Maiani et al. (2005, 2019b)
@@
$H_{ss}=\sum_{i<j}2\kappa_{ij}(\mathbf{s}_{i}\cdot\mathbf{s}_{j}),$ (3)
where $m_{i}$ represents the constituent mass of each quark, $E$ is the energy eigenvalue of the tetraquark system obtained based on the BO approximation, and $H_{ss}$ corresponds to the color-spin hyperfine interaction.
II.2 Born-Oppenheimer potention
@@
$\begin{split}\left(\sum_{\text{light}}\frac{p_{i}^{2}}{2m}\right)f&+V_{I}(x_{A,B},x_{1,2})f\\ &=\mathcal{E}(x_{A,B})f,\end{split}$ (4)
where $f=f(x_{A,B},x_{1,2})$ is the eigenfunction for the light quarks. $\mathcal{E}(x_{A,B})$ is the eigenvalue dependent on the positions of the heavy quarks $x_{A,B}$. $V_{I}(x_{A,B},x_{1,2})$ describes the interaction between the light and heavy quarks. The total wavefunction can be written as
$\Psi=\psi(x_{A,B})f(x_{A,B},x_{1,2}),$ (5)
where $\psi$ describes the heavy quarks, and $f$ describes the light quarks. Under the Born-Oppenheimer approximation, we apply $H$ to $\Psi$ , assuming that the wavefunction $f$ for the light quarks is sensitive to changes in the positions of the heavy quarks, so $\nabla_{i}f$ cannot be neglected. But the wavefunction $\psi$ for the heavy quarks changes slowly with respect to the positions of the heavy quarks, so $\nabla_{i}\psi$ can be approximated as zero.
Thus, for the kinetic term
$-\hbar^{2}\nabla_{i}^{2}\Psi=-\hbar^{2}\left(f\nabla_{i}^{2}\psi+2(\nabla_{i}\psi)\cdot(\nabla_{i}f)+\psi\nabla_{i}^{2}f\right),$ (6)
it can be approximated as
$-\hbar^{2}\nabla_{i}^{2}\Psi\approx-\hbar^{2}\left(f\nabla_{i}^{2}\psi+\psi\nabla_{i}^{2}f\right).$ (7)
We can rewrite the Schrödinger equation as
$\left(\sum_{\text{heavy}}\frac{p_{i}^{2}}{2M}+V_{\text{BO}}(x_{A,B})\right)\psi=E\psi,$ (8)
@@
$V_{\text{BO}}(x_{A,B})=V(x_{A,B})+\mathcal{E}(x_{A,B})+V_{\text{conf}}.$ (9)
Next, we will conduct a detailed analysis of the specific form of $V_{\text{BO}}$. We recall that the color coupling between any pair of particles in color representation $R$ is given by
$V_{C}(r)=\lambda_{q_{1}q_{2}}(\mathbf{R})\frac{\alpha_{s}}{r},$ (10)
where $q_{1}$ and $q_{2}$ represent the irreducible representations of the quark pair.
@@
For hidden heavy flavor tetraquarks, the color of the two heavy quarks after coupling can be either $1$ or $8$. In the first case, however, the interaction between $Q\bar{Q}$ and $q\bar{q}$ pairs occurs through the exchange of color singlets. This leads to a model dominated by nuclear-like forces, ultimately resulting in the formation of hadrocharmonium Dubynskiy and Voloshin (2008). Therefore, in our calculation, we exclude the scenario where $Q\bar{Q}$ couples in the color state $1$. So $\lambda_{Q\bar{Q}}$ and $\lambda_{q\bar{q}}$ are given by
$\lambda_{q_{1}q_{2}}(\mathbf{R})=\frac{1}{2}\left[C_{2}(\mathbf{R})-C_{2}(q_{1})-C_{2}(q_{2})\right],$ (11)
where $C_{2}$ is the second Casimir operator. We note the following results:
$C_{2}(1)$ $=0,\quad C_{2}(\mathbf{R})=C_{2}(\mathbf{\bar{R}}),$ (12)
$C_{2}(3)$ $=\frac{4}{3},\quad C_{2}(6)=\frac{10}{3},\quad C_{2}(8)=3.$
@@
In order to derive $\mathcal{E}(x_{A,B})$, we need to solve Schrödinger Eq.(4). Given that the impact of heavy quarks on light quarks is non-negligible, We adopt a method similar to the treatment of hydrogen molecules to handle $Q\bar{Q}q\bar{q}$. Specifically, we model the hidden heavy flavor tetraquarks as two combined hydrogen-like atoms. In this model, a light particle is paired exclusively with one heavy particle, forming two symmetrical orbits. The total energy of the two orbits is $E_{0}$, and the interactions not included in the orbits are treated as perturbations, denoted as $\Delta E$. This gives rise to two possible cases, $Qq$ and $\bar{Q}\bar{q}$, and $\bar{Q}q$ and $Q\bar{q}$. The second case leads to a molecular state of two meson-like combinations, which is not considered in our analysis. We focus only on the tightly bound states formed by the $Qq$ and $\bar{Q}\bar{q}$ combinations.To analyze the interaction between the orbit $Qq$, we introduce the Cornell potential, a prototype for non-relativistic quark interactions.
$V(r)=-\frac{4}{3}\frac{\alpha_{s}}{r}+\sigma r+V_{0},$ (13)
where $r$ is the effective radius of the hadron, $\alpha_{s}$ is the QCD coupling constant, $\sigma$ is the QCD string tension, and $V_{0}\simeq-0.3\,\text{GeV}$ is a constant.
But the range of the Cornell potential is limited to heavy quarkonium or interactions between quarks with similar masses. For other cases, such as $qQ$, we modify the formula as follows
$V_{q_{1}q_{2}(\bar{q})}(r)=\lambda_{q_{1}q_{2}}\frac{\alpha_{S}}{r}+\frac{3|\lambda_{q_{1}q_{2}}|}{4}\sigma r+V_{0},$ (14)
where $\sigma$ is the string tension extracted from the charmonium spectrum. As for $\lambda_{q_{1}q_{2}}$, in the tetraquark system, $q_{1}$ and $q_{2}$ are in a superposition of two $SU(3)_{c}$ representations with respective amplitudes $a$ and $b$. then we have
$T=a\left[(q_{1}q_{2})_{\mathbf{R}_{1}}\cdots\right]_{\mathbf{R}_{1}}+b\left[(q_{1}q_{2})_{\mathbf{R}_{2}}\cdots\right]_{\mathbf{R}_{2}},$ (15)
$\lambda_{q_{1}q_{2}}=a^{2}\lambda_{q_{1}q_{2}}(\mathbf{R}_{1})+b^{2}\lambda_{q_{1}q_{2}}(\mathbf{R}_{2}).$ (16)
For the $Qq$ orbit, We use the expression from (14) to represent the potential between the $Qq$ quarks,
$V_{Qq}=-\frac{1}{3}\frac{\alpha_{S}}{r}+\frac{1}{4}\sigma+V_{0}.$ (17)
@@
$R(r)=\frac{A^{3/2}}{\sqrt{\pi}}e^{-Ar}.$ (18)
Substituting the model parameters, we can obtain the ground state and the corresponding wave function by using variational calculation.
Due to the high symmetry of the tetraquark system, the wave functions of the two non-interacting orbits can be expressed as
$f(1,2)=\psi(1)\phi(2)=R(|x_{1}-x_{A}|)R(|x_{2}-x_{B}|).$ (19)
The ground state energy corresponding to the wave function is
$E_{0}=2(\langle H\rangle_{\min}+V_{0}).$ (20)
Next, we calculate the perturbation term as
$\begin{split}H_{\text{pert}}=&-\frac{7}{6}\alpha_{S}\left(\frac{1}{|x_{1}-x_{B}|}+\frac{1}{|x_{2}-x_{A}|}\right)\\ &+\frac{1}{6}\alpha_{S}\frac{1}{|x_{1}-x_{2}|}.\end{split}$ (21)
The perturbative energy is given by
$\Delta E=\langle f|H_{\text{pert}}|f\rangle,$ (22)
Further calculation yields
$\Delta E=-\frac{7}{6}\alpha_{S}2I_{1}(r_{AB})+\frac{1}{6}\alpha_{S}I_{4}(r_{AB}),$ (23)
where the functions $I_{1}(r)$ and $I_{4}(r)$ are given by
$I_{1}(r_{AB})=\int d^{3}\xi|\psi(\xi)|^{2}\frac{1}{|\xi-x_{B}|},$ (24)
$I_{4}(r_{AB})=\int d^{3}\xi d^{3}\eta|\psi(\xi)|^{2}|\phi(\eta)|^{2}\frac{1}{|\xi-\eta|}.$ (25)
We obtain the final energy as
$\mathcal{E}(x_{A,B})=E_{0}+\Delta E(x_{A,B}).$ (26)
Finally, we consider the color confinement of the entire tetraquark system. Recall that $3\otimes 3=\bar{3}\oplus 6$, and we observe that the $cq$ orbit is not a singlet, implying it is affected by the confinement potential. We need to include a linear string tension to simulate the confinement effect, which is given by
$V_{\text{conf}}(r)=\sigma_{T}\times(r-R_{0})\times\theta(r-R_{0}).$ (27)
$R_{0}$ is the relative distance between quarks within a hadron. In principle, $R_{0}$ should be treated as a free parameter, and its value can be determined by experimental phenomena related to the tetraquark states. For parameter $\sigma_{T}$, we use the Fierz identity to express the tetraquark state $T=|(c\bar{c})_{\bm{8}}(\bar{q}q)_{\bm{8}}\rangle_{\bm{1}}$ as
$T=\sqrt{\frac{2}{3}}|(Qq)_{\bar{\bm{3}}}(\bar{Q}\bar{q})_{\bm{3}}\rangle_{\bm{1}}-\sqrt{\frac{1}{3}}|(Qq)_{\bm{6}}(\bar{Q}\bar{q})_{\bar{\bm{6}}}\rangle_{\bm{1}}.$ (28)
Due to the gluon screening effect, the diquark in the six-dimensional representation ($6$) adds a component to the three-dimensional representation ($\bar{3}$), making $\sigma_{T}$ closer to $\sigma$. For simplicity, we adopt $\sigma_{T}=\sigma$. Thus, the final potential is
$\begin{split}V_{\text{BO}}(r)=&\frac{1}{6}\alpha_{S}\frac{1}{r_{AB}}+E_{0}+\Delta E(x_{A,B})\\ &+\sigma\times(r-R_{0})\times\theta(r-R_{0}).\end{split}$ (29)

### v2

@@ Abstract
The Born-Oppenheimer approximation is one of the very successful tools for solving the hydrogen atom problem. The experimental discovery of hidden heavy flavor tetraquarks, $Q\bar{Q}q\bar{q}$ ($Q=c,b$ and $q=u,d,s$), provides great possibilities for the hydrogen-bond-like structure of the Quantum Chromodynamics version. In this work, considering that the colors of $Q\bar{Q}$ and $q\bar{q}$ are both $8$, the tetraquark $Q\bar{Q}q\bar{q}$ system is formed by color coupling $8\otimes 8\rightarrow 1$. In order to study the mass splitting caused by the color-spin hyperfine interaction, the color-spin basis vectors of the $S$-wave tetraquark states are appropriately constructed. Then we use the Born-Oppenheimer approximation to calculate the mass spectra of the $S$-wave hidden heavy flavor tetraquark states. The results show that some of the hidden heavy flavor exotic hadrons discovered experimentally can be well explained as this type of hydrogen-bond-like tetraquark structure. In addition, some candidates for tetraquark bound states are predicted and may be compact tetraquark states.
@@ I Introduction
Quantum chromodynamics (QCD) provides a rich framework for understanding hadrons, extending beyond the conventional mesons (one quark, one antiquark) and baryons (three quarks) to encompass exotic configurations such as multiquark states, hybrids, and glueballs Godfrey and Isgur (1985); Klempt and Zaitsev (2007); Lebed et al. (2017). Among the multiquark states, the study of tetraquark states (composed of two quarks and two antiquarks) have received significant attention Jaffe (1977a, b); Lodha and Rai (2024). Notably, certain hidden heavy flavor tetraquarks, $Q\bar{Q}q\bar{q}$ ($Q=c,b$ and $q=u,d,s$), can be interpreted as analogues of the hydrogen bonds within QCD version. This allows the interaction between heavy quark-antiquark pairs and light quarks can be treated using the Born-Oppenheimer (BO) approximation Esposito et al. (2013); Braaten et al. (2014); Maiani et al. (2019a, b); Mutuk (2024); Berwein et al. (2024, 2024); Germani et al. (2025). Based on the fact that heavy quarks are much heavier than light quarks, we can decouple the heavy quark dynamics from the light quark dynamics. This strategy makes it possible to systematically study the mass spectra and structures of the hidden heavy flavor tetraquark states.
@@
Recently, there is increasing experimental evidence proving the existence of possible hidden heavy flavor tetraquark states. In 2020, the first candidate for a charged hidden-charm tetraquark state with strangeness, $Z_{cs}(3985)^{-}$, was discovered in the processes of $e^{+}e^{-}\rightarrow K^{+}D_{s}^{-}D^{\ast 0}$ and $K^{+}D_{s}^{\ast-}D^{0}$ by the BESIII collaboration Ablikim et al. (2021). Its non-zero charge characteristic enables it to be clearly distinguished from hybrid or charmonium. Theoretically, $Z_{cs}(3985)^{-}$ is generally interpreted as tetraquark state $cs\bar{c}\bar{u}$ with a quantum number assignment $J^{P}=1^{+}$ Meng et al. (2020); Karliner and Rosner (2021); Giron et al. (2021); Guo et al. (2023); Liu et al. (2025). Later in 2021 the LHCb collaboration reported two hidden-charm exotic structures in the decay $B^{+}\rightarrow J/\psi\phi K^{+}$, referred to as $Z_{cs}(4000)^{+}$ and $Z_{cs}(4220)^{+}$ Aaij et al. (2021). Their quark composition is thought to be $cu\bar{c}\bar{s}$, and their spin-parity quantum quantum numbers are determined to be $J^{P}=1^{+}$ and $1^{+}$ (or $1^{-}$), respectively Aaij et al. (2021). However, theoretical studies suggest that the $Z_{cs}(4000)^{+}$ and $Z_{cs}(4220)^{+}$ may be candidates for charmonia A. and Dhir (2024); Bokade and Bhaghyesh (2025), tetraquarks Li et al. (2024); Vogt (2024); Yu et al. (2024), or molecules Yang et al. (2021); Chen (2024). It is worth noting that the BESIII collaboration recently observed a new $Z_{cs}(4123)^{-}$ in the processes of $e^{+}e^{-}\rightarrow K^{+}D_{s}^{\ast-}D^{\ast 0}+c.c.^{\ast}$, and its quantum number assignment may be $J^{P}=1^{+}$ Ablikim et al. (2023). The discovery of the $Z_{cs}(4123)^{-}$ may add a new member to the family of charged hidden-charm tetraquark with strangeness. In 2023, A near-threshold peaking structure $X(3960)$ is observed in the $D_{s}^{+}D_{s}^{-}$ invariant mass spectrum by the LHCb collaboration, and the quantum number assignment $J^{PC}=0^{++}$ is favored Aaij et al. (2023). Research shows that the nature of $X(3960)$ may also be a tetraquark $cs\bar{c}\bar{s}$ Badalian and Simonov (2023); Hoffer et al. (2024); Chen et al. (2025). Obviously, these exotic hadrons all have the common components $Q\bar{Q}q\bar{q}$, and figuring out their nature will help us better understand QCD.
@@
In this work, we apply the BO approximation to calculate the mass spectra of tetraquark $Q\bar{Q}q\bar{q}$. The heavy $Q\bar{Q}$ forms a color octet, which combines with a light $q\bar{q}$ octet to produce a color-singlet tetraquark, that is, the color coupling is represented as $8\otimes 8\rightarrow 1$. Given the same amount of momentum, the heavy quarks move much more slowly than the light quarks. The light quark (or antiquark) interactions can be treated as perturbations to the dominant heavy-quark potential, allowing the Schrödinger equation for the heavy quark separation to be solved numerically. The calculation of diquark-antidiquark paradigm was extended to tetraquark system. Using $X(3872)$ as input, other possible tetraquark states were studied.
@@
This paper is organized as follows. In Sec. II, We give a brief introduction of the tetraquark Hamiltonian based on the BO approximation. We obtain the BO potential using the perturbation method and construct the basis vector wave functions of the tetraquark system $Q\bar{Q}q\bar{q}$. In Sec. III, we determine the model parameters, such as the effective masses of quarks and the color coupling strengths. The mass spectra and decay modes of various hidden heavy tetraquark states are calculated and discussed in this section. We summarize in Sec. IV. The nature units $c=\hbar=1$ are used throughout.
@@ II.1 Hamiltonian in the Born-Oppenheimer approximation
${\cal H}$ $=$ $\sum_{i=1}^{4}\left(\frac{p_{i}^{2}}{2m_{i}}\right)+V(\bm{x}_{A,B})$ (1)
$+$ $V_{I}(\bm{x}_{A,B},\bm{x}_{1,2})+V_{\text{conf}},$
where $\bm{x}_{A,B}$ and $\bm{x}_{1,2}$ represent the coordinates of heavy quarks and light quarks respectively, $m_{i}$ represents the effective mass of the $i$-th constituent quark. $\sum_{i=1}^{4}\left(\frac{p_{i}^{2}}{2m_{i}}\right)$ is defined as the kinetic energy of the quarks. $V(\bm{x}_{A,B})$ describes the interaction between the heavy quarks. $V_{I}(\bm{x}_{A,B},\bm{x}_{1,2})$ describes the general interactions involving light-heavy and light-light quarks. $V_{\text{conf}}$ is defined as the color confinement part of the tetraquark states.
If we focus only on the $S$-wave tetraquark states, the spatial degrees of freedom can be encoded into the model parameters. Considering the energy level splitting caused by the color-spin interaction, the Hamiltonian for the $S$-wave tetraquark states be simplified as Maiani et al. (2005, 2019b)
@@
with
$H_{ss}=\sum_{i<j}2\kappa_{ij}(\bm{s}_{i}\cdot\bm{s}_{j}),$ (3)
where $E$ is the energy eigenvalue of the tetraquark system obtained based on the BO approximation. The coefficient $\kappa_{ij}$ is inversely proportional to the product of the masses of the $i$-th and $j$-th constituent quarks, and directly proportional to the color coupling strength of the corresponding quark pair. The Hamiltonian $H_{ss}$ is the color-spin hyperfine interaction.
II.2 Born-Oppenheimer potential
@@
$\begin{split}\left[\left(\sum_{\text{light}}\frac{p_{i}^{2}}{2m}\right)+V_{I}(\bm{x}_{A,B},\bm{x}_{1,2})\right]f=\mathcal{E}(\bm{x}_{A,B})f,\end{split}$ (4)
where $f=f(\bm{x}_{A,B},\bm{x}_{1,2})$ is the eigenfunction for the light quarks. $\mathcal{E}(\bm{x}_{A,B})$ is the eigenvalue dependent on the positions of the heavy quarks. The total wavefunction of tetraquark system can be written as
$\Psi=\psi(\bm{x}_{A,B})f(\bm{x}_{A,B},\bm{x}_{1,2}),$ (5)
where $\psi$ describes the wavefunction of heavy quarks. We let the Hamiltonian Eq.(1) act on the wave function Eq.(5). Since the wavefunction $f$ of light quarks is very sensitive to changes in the position of heavy quarks, $\nabla_{i}f$ cannot be ignored. But the wavefunction $\psi$ of heavy quarks changes slowly with respect to the positions of the heavy quarks, $\nabla_{i}\psi$ can be approximated as zero. For the kinetic term, it can be approximated as
$\nabla_{i}^{2}\Psi\approx\left(f\nabla_{i}^{2}\psi+\psi\nabla_{i}^{2}f\right).$ (6)
Now we can rewrite the Schrödinger equation of heavy quarks as
$\left(\sum_{\text{heavy}}\frac{p_{i}^{2}}{2M}+V_{\text{BO}}(\bm{x}_{A,B})\right)\psi=E\psi,$ (7)
@@
$V_{\text{BO}}(\bm{x}_{A,B})=V(\bm{x}_{A,B})+\mathcal{E}(\bm{x}_{A,B})+V_{\text{conf}}.$ (8)
Next, we analyze the specific form of $V_{\text{BO}}$ potential. In a hadron system, the color coupling between any quark and quark (or antiquark) can be expressed as
$V(r)=\lambda_{q_{1}q_{2}}(\bm{R})\frac{\alpha_{s}}{r},$ (9)
with the Casimir operator
$\lambda_{q_{1}q_{2}}(\bm{R})=\frac{1}{2}\left[C(\bm{R})-C(q_{1})-C(q_{2})\right],$ (10)
where $r$ is the relative distance between the two quarks, and $\alpha_{s}$ is the QCD running coupling constant. The notation $\bm{R}$ is the color representation of any pair of particles, $q_{1}$ and $q_{2}$ represent the irreducible representations of the quark. In the SU(3) group representation, we can get $C(3)=4/3$, $C(6)=10/3$, $C(1)=0$, $C(8)=3$, $C(\bar{q})=C(q)$, and $C(\bm{\bar{R}})=C(\bm{R})$.
@@
For hidden heavy flavor tetraquarks, the color couplings of the two heavy quarks can be either $1$ or $8$. In the first case, however, the interaction between $Q\bar{Q}$ and $q\bar{q}$ pairs occurs through the exchange of color singlets. This is similar to the nucleon-nucleon interaction, which manifests itself as the configuration of hadronic molecular states. Therefore,we exclude the scenario where $Q\bar{Q}$ couples in the color state $1$. In our calculation, we only consider the coupling of quark pairs both ${Q\bar{Q}}$ and ${q\bar{q}}$ with color state $8$.
@@
In order to derive $\mathcal{E}$, we need to solve Schrödinger Eq.(4). Given that the impact of heavy quarks on light quarks is non-negligible, the hidden heavy flavour tetraquark state can be considered as a hydrogen-like molecular system of QCD version. Then, we can extend the calculation methods for hydrogen molecules to study the tetraquark $Q\bar{Q}q\bar{q}$. Specifically, a light quark is paired exclusively with one heavy quark (or antiquark) forming two symmetrical orbits. To analyze the interaction between the orbit $\bar{Q}q$ or $Q\bar{q}$, we introduce the Cornell potential of non-relativistic quark model,
$V_{\bar{Q}q}(r)=-\frac{4}{3}\frac{\alpha_{s}}{r}+\sigma r+V_{0},$ (11)
where $\sigma$ is the string tensor strength of QCD and is proportional to $|\lambda|$. The parameter $V_{0}\simeq-0.315\,\text{GeV}$ can be obtained by fitting the quarkonium. For the orbit $Qq$ or $\bar{Q}\bar{q}$, we modify the interaction as follows
$V_{Qq}(r)=-\frac{1}{3}\frac{\alpha_{s}}{r}+\frac{1}{4}\sigma r+V_{0}.$ (12)
@@
$R(r)=\frac{A^{3/2}}{\sqrt{4\pi}}e^{-Ar},$ (13)
where $A$ is the variational parameter. Substituting the trial wave function and interaction into the Hamiltonian, we can obtain the ground state energy ($\langle H\rangle_{\text{min}}$) and the corresponding wave function ($\zeta$ or $\xi$) of a heavy-light quark pair by using variational calculation.
We consider as nonperturbative process for the orbits of two non-interacting heavy-light quark pairs. Due to the symmetry of the tetraquark system, the wave function of light quarks can be expressed as
$f=\zeta(\bm{\rho})\xi(\bm{\eta})=R(|\bm{x}_{1}-\bm{x}_{A}|)R(|\bm{x}_{2}-\bm{x}_{B}|),$ (14)
and the ground state energy corresponding to the wave function reads
$E_{Qq\leftrightarrow\bar{Q}\bar{q}}=2\langle H\rangle_{\text{min}}.$ (15)
For the orbit formed by the light quark-antiquark pair, we regard it as the perturbation term. Then, the perturbation Hamiltonian is
$H_{p}=$displaystyle-- $\frac{7}{6}\alpha_{s}\left(\frac{1}{|\bm{x}_{1}-\bm{x}_{B}|}+\frac{1}{|\bm{x}_{2}-\bm{x}_{A}|}\right)$ (16)
$+$ $\frac{1}{6}\alpha_{s}\frac{1}{|\bm{x}_{1}-\bm{x}_{2}|}.$
For simplicity, the perturbative energy is written by
$\Delta E=\langle f|H_{p}|f\rangle,$ (17)
By calculation, $\Delta E$ is written as
$\Delta E=-\frac{7}{6}\alpha_{s}2I_{1}(r_{AB})+\frac{1}{6}\alpha_{s}I_{2}(r_{AB}),$ (18)
where $r_{AB}=|\bm{x}_{B}-\bm{x}_{A}|$. Functions $I_{1}$ and $I_{2}$ are
$I_{1}(r_{AB})=\int\text{d}^{3}\rho|\psi(\rho)|^{2}\frac{1}{|\bm{\rho}-\bm{r}_{AB}|},$ (19)
$I_{2}(r_{AB})=\int\text{d}^{3}\rho\,\text{d}^{3}\eta|\psi(\rho)|^{2}|\phi(\eta)|^{2}\frac{1}{|\bm{\rho}-\bm{\eta}|}.$ (20)
Now, we can obtain the final energy as
$\mathcal{E}(\bm{x}_{A,B})=E_{Qq\leftrightarrow\bar{Q}\bar{q}}+\Delta E.$ (21)
For the hidden heavy flavor tetraquark system, the overall color coupling needs to ensure color singlet. We need to include a linear string tension to simulate the confinement effectBruschini and González (2020); Andreev (2022). Considering the tetraquark configuration $|(Q\bar{Q})_{\bm{8}}(q\bar{q})_{\bm{8}}\rangle_{\bm{1}}$, the linearly increasing confinement term reads Maiani et al. (2019b)
$V_{\text{conf}}=\sigma\times(r-R_{0})\times\theta(r-R_{0}).$ (22)
$R_{0}$ is the relative distance between quarks within a hadron. In general, $R_{0}$ can be treated as an adjustable free parameter and can be obtained by fitting the hadron spectrum. Therefore, the Born-Oppenheimer potential can be expressed as
$V_{\text{BO}}(r_{AB})$ $=$ $\frac{1}{6}\alpha_{S}\frac{1}{r_{AB}}+E_{Qq\leftrightarrow\bar{Q}\bar{q}}+\Delta E$ (23)
$+$ $\sigma\times(r_{AB}-R_{0})\times\theta(r_{AB}-R_{0}).$
Substituting Eq.(23) into Eq.(7), we can obtain the binding energy $E$ of the hidden heavy flavor tetraquark state under the BO approximation.